# A Federated Distributionally Robust Support Vector Machine with Mixture of Wasserstein Balls Ambiguity Set for Distributed Fault Diagnosis

### v1

@@ Abstract
The training of classification models for fault diagnosis tasks using geographically dispersed data is a crucial task for original parts manufacturers (OEMs) seeking to provide long-term service contracts (LTSCs) to their customers. Due to privacy and bandwidth constraints, such models must be trained in a federated fashion. Moreover, due to harsh industrial settings the data often suffers from feature and label uncertainty. Therefore, we study the problem of training a distributionally robust (DR) support vector machine (SVM) in a federated fashion over a network comprised of a central server and $G$ clients without sharing data. We consider the setting where the local data of each client $g$ is sampled from a unique true distribution $\mathbb{P}_{g}$, and the clients can only communicate with the central server. We propose a novel Mixture of Wasserstein Balls (MoWB) ambiguity set that relies on local Wasserstein balls centered at the empirical distribution of the data at each client. We study theoretical aspects of the proposed ambiguity set, deriving its out-of-sample performance guarantees and demonstrating that it naturally allows for the separability of the DR problem. Subsequently, we propose two distributed optimization algorithms for training the global FDR-SVM: i) a subgradient method-based algorithm, and ii) an alternating direction method of multipliers (ADMM)-based algorithm. We derive the optimization problems to be solved by each client and provide closed-form expressions for the computations performed by the central server during each iteration for both algorithms. Finally, we thoroughly examine the performance of the proposed algorithms in a series of numerical experiments utilizing both simulation data and popular real-world datasets.
@@ 1 Introduction
Original Equipment Manufacturers (OEMs) of capital-intensive assets such as gas turbines, aircraft engines, and mining equipment are increasingly implementing long-term service contracts (LTSCs) that can last for 10 to 20 years to ensure their products perform reliably over time [1]. Such contracts are beneficial to both the OEMs and the customers, as they provide the OEMs with a steady revenue stream, whereas they relieve the customers from all diagnosis and repair tasks in the occurrence of fault events. Indeed, one of the key services provided by the OEM as part of LTSCs is the diagnosis of any faults that may occur in the equipment used by the customers. Recently, many research efforts have been dedicated to performing fault diagnosis tasks by leveraging a wide array of data-driven Machine Learning (ML) classifiers. For example, ML models have been used in the fault diagnosis of motors [2], rotary machines [3], and pumps [4].
@@
    We propose the MoWB ambiguity set, which extends the conventional Wasserstein ball to the distributed setting. Our proposed ambiguity set contains mixtures of distributions from Wasserstein balls defined locally at each client. Using this ambiguity set, we reformulate the FDR-SVM problem into a separable form that can be solved in a distributed manner without sharing data. We also derive out-of-sample performance guarantees for the MoWB ambiguity set under certain assumptions.
@@
The rest of this paper is organized as follows. Section 2 provides a summary of the works relevant to the FDR-SVM, covering topics of DRO, FL, and their combination. Section 3 provides a brief background on the theoretical aspects of Wasserstein DRO, as well as a summary of the problem setup. Section 4 introduces the proposed MoWB ambiguity set and demonstrates how the MoWB ambiguity set leads to problem separability. Section 5 describes the algorithms proposed to train the FDR-SVM and discusses their theoretical guarantees. Section 6 contains our numerical experiments, and finally Section 7 provides some concluding remarks.
@@ 2.1 Distributionally Robust Classifiers
Many existing works in the literature seek to develop DR versions of common classifiers to improve their out-of-sample performance. For example, the authors in [7] and [15] develop Wasserstein DR versions of logistic regression and hard and soft-margin SVMs, respectively. They provide tractable, convex reformulations for their proposed DR problems, leveraging popular duality results from [10, 9]. This work is extended in [23], where the authors derive a tractable, convex reformulation of a Wasserstein DR logistic regression model for data with mixed features, demonstrating that considering categorical data leads to a fundamentally different problem.
@@
Finally, the authors in [20] utilize group DRO to address the tendency of classification deep neural networks (DNNs) to learn spurious correlations in the data. More specifically, they manually group the training data according to those correlations and minimize the risk with respect to the worst-case group. However, they key challenge in that work is that data grouping has to be performed manually, which can be extremely difficult for very large datasets. This challenge is addressed by the authors in [25], where they utilize a DNN to perform the data grouping. The common aspect shared between the previously listed works is that they rely on the training data being present at a central location. Indeed as argued in [18], this can make such models difficult to extend to distributed settings.
@@ 2.3 Distributionally Robust Federated Learning
Another research direction that combines elements of DRO and FL is explored in [37] where the authors develop a Wasserstein ambiguity set using barycenters. In this setting, the authors consider multiple clients, each of which has local samples from $\mathbb{P}$ allowing it to construct a local empirical distribution ^$\widehat{\mathbb{P}}_{N_{g}}$. The global ambiguity set is then centered around the barycenter (which is similar to a mean) of all {^$\{\widehat{\mathbb{P}}_{N_{g}}\}_{g=1}^{G}$. However, the barycenter is not always guaranteed to exist and can be difficult to compute in a distributed fashion when it does. Similarly, the authors in [18] study a multi-client network where each client has access to a finite number of samples from $\mathbb{P}$. The clients seek to collaborate via P2P communication to solve a Wasserstein DRO problem in a distributed fashion. The authors show that the reformulation of the Wasserstein DRO problem in [10, 9] admits a reformulation that can be solved in a distributed fashion via Lagrange duality if the clients can communicate with each other.
@@ 4.1 Problem Separability
In this section, we extend the traditional Wasserstein ball to the distributed setting via the novel MoWB ambiguity set $\mathcal{A}_{G}$ defined next.
@@
Observe that when $G=1$, our proposed ambiguity set reduces to the conventional Wasserstein ball $\mathcal{A}_{\varepsilon,q,d}(\Xi)$ defined in (2). This suggests that our proposed ambiguity set indeed naturally extends the conventional Wasserstein ball to the distributed setting, allowing for the problem to be reformulated in a separable fashion as we demonstrate next.
@@
$\inf_{\boldsymbol{w}}\sup_{\mathbb{Q}\in\mathcal{A}_{G}}\mathbb{E}^{\mathbb{Q}}[\ell_{H}(\boldsymbol{w};\boldsymbol{\xi})]$ (6a)
$=\inf_{\boldsymbol{w}}\sup_{\left\{\mathbb{Q}_{g}\in\mathcal{A}^{(g)}_{\varepsilon_{g},1,d}(\Xi)\right\}_{g=1}^{G}}\mathbb{E}^{\sum_{g=1}^{G}\alpha_{g}\mathbb{Q}_{g}}[\ell_{H}(\boldsymbol{w};\boldsymbol{\xi})]$ (6b)
$=\inf_{\boldsymbol{w}}\sup_{\left\{\mathbb{Q}_{g}\in\mathcal{A}^{(g)}_{\varepsilon_{g},1,d}(\Xi)\right\}_{g=1}^{G}}\sum_{g=1}^{G}\alpha_{g}\mathbb{E}^{\mathbb{Q}_{g}}[\ell_{H}(\boldsymbol{w};\boldsymbol{\xi})]$ (6c)
$=\inf_{\boldsymbol{w}}\sum_{g=1}^{G}\alpha_{g}\sup_{\mathbb{Q}_{g}\in\mathcal{A}^{(g)}_{\varepsilon_{g},1,d}(\Xi)}\mathbb{E}^{\mathbb{Q}_{g}}[\ell_{H}(\boldsymbol{w};\boldsymbol{\xi})],$ (6d)
where (6b) follows from the definition of the global ambiguity $\mathcal{A}_{G}$ set in (4), (6c) follows from the law of expectation over mixture distributions, and (6d) follows by recognizing that the maximization problems are separable due to each decision variable only affecting its corresponding term. ∎
@@ 4.2 Out-of-Sample Performance Guarantees
All the training samples $\{(\widehat{\boldsymbol{x}}_{(n_{g})},\widehat{y}_{(n_{g})})\}_{n=1}^{N}$ at client $g$ are IID samples from its true distribution $\mathbb{P}_{g}$.
@@
Beyond its desirable out-of-sample performance properties, our proposed MoWB ambiguity set also inherits other desirable properties from the traditional Wasserstein ambiguity set listed in [9]. One key property is the ability of the MoWB ambiguity set to anticipate black swans by possibly assigning probability mass to any point in the support set $\Xi$ regardless of the structure of ^$\widehat{\mathbb{P}}_{N_{g}}$ for all $g\in[G]$.
@@ 5.1 SM Algorithm
$\max_{\begin{subarray}{c}\beta_{n_{g}}^{+},\beta_{n_{g}}^{-}\\ \boldsymbol{q}_{n_{g}}^{+},\boldsymbol{q}_{n_{g}}^{-}\end{subarray}}H_{g}(\boldsymbol{w})=\left\{\begin{aligned} &\max_{\begin{subarray}{c}\beta_{n_{g}}^{+},\beta_{n_{g}}^{-}\\ \boldsymbol{q}_{n_{g}}^{+},\boldsymbol{q}_{n_{g}}^{-}\end{subarray}}&&\frac{1}{N_{g}}\sum_{n_{g}=1}^{N_{g}}\big{(}-(\beta_{n_{g}}^{+}-\beta_{n_{g}}^{-})\widehat{y}_{n_{g}}\boldsymbol{w}^{\mathsf{T}}\widehat{\boldsymbol{x}}_{n_{g}}-\widehat{y}_{n_{g}}\boldsymbol{w}^{\mathsf{T}}(\boldsymbol{q}_{n_{g}}^{+}-\boldsymbol{q}_{n_{g}}^{-})\big{)}&&\\ &\operatorname{s.t.}&&\sum_{n_{g}=1}^{N_{g}}\big{(}||\boldsymbol{q}_{n_{g}}^{+}||+||\boldsymbol{q}_{n_{g}}^{-}||+\kappa_{g}\beta_{n_{g}}^{-}\big{)}\leq N_{g}\varepsilon_{g}\\ &&&\beta_{n_{g}}^{+}+\beta_{n_{g}}^{-}=1&&\forall n_{g}\in[N_{g}]\\ &&&0\leq\beta_{n_{g}}^{+}\widehat{\boldsymbol{x}}_{n_{g}}-\boldsymbol{q}_{n_{g}}^{+}\leq\beta_{n_{g}}^{+}&&\forall n_{g}\in[N_{g}]\\ &&&0\leq\beta_{n_{g}}^{-}\widehat{\boldsymbol{x}}_{n_{g}}-\boldsymbol{q}_{n_{g}}^{-}\leq\beta_{n_{g}}^{-}&&\forall n_{g}\in[N_{g}]\\ &&&\beta_{n_{g}}^{+},\beta_{n_{g}}^{-}\geq 0&&\forall n_{g}\in[N_{g}]\end{aligned}\right.,$ (9)
@@
The computed local subgradients are sent to the central server, which computes a weighted sum of the subgradients using weights $\alpha_{g}$, and performs a subgradient step on the model parameters $\boldsymbol{w}$. The central server then sends the updated $\boldsymbol{w}$ to the clients and the process repeats until convergence, which we discuss in further detail later. We write our SM algorithm explicitly as follows:
@@ 5.2 ADMM Algorithm
The ADMM algorithm requires each client $g$ to solve their own local problem and shares their optimal local model $\boldsymbol{w}^{\ast}_{g}$ with the central server. The central server then proceeds to aggregate the local models to obtain an optimal global model $\boldsymbol{w}^{\ast}$, which it sends to the clients. The process repeats until convergence, which we discuss in further detail later.
@@
[…]
@@ 5.3.1 SM Algorithm
The SM Algorithm 1 converges to an optimal solution $\boldsymbol{w}^{\ast}$ of the problem in (5) with arbitrary tolerance $\epsilon_{1}$ if the stepsize $\gamma(t)\rightarrow 0$ as $t\rightarrow\infty$ and $\sum_{t=1}^{\infty}\gamma(t)=\infty$.
@@
Note that it is established in [42] that the subgradient method is guaranteed convergence if three conditions hold: i) the objective function is convex, ii) the objective function is Lipschitz continuous, and iii) the step-size is diminishing an appropriate rate as stated in the Theorem statement. We verify the first two conditions in Lemmas 2 and 3, whereas the third condition can be ensured by choosing an appropriate step size. ∎
@@
Suppose that the $\ell_{\infty}$-norm is used in the worst-case distribution problem (9), and that the barrier method equipped with the log barrier function and Newton updates is used to solve it. Then, the SM algorithm 1 equipped with an appropriately diminishing step-size [42] has a worst-case time complexity of $\mathcal{O}\left(\epsilon_{1}^{-2}\left[N_{g^{\ast}}^{3.5}P^{3.5}\log(\epsilon_{2}^{-1})+GP\right]\right)$, where $\epsilon_{1}$ and $\epsilon_{2}$ are the tolerances of the solutions of the subgradient method and the worst-case distribution problem (9), respectively, and $N_{g^{\ast}}$ is the greatest number of samples available at any one client. This analysis can be directly extended to cases where other norms are used in (9).
@@
We begin by examining the time complexity of problem (9) that each client $g$ solves at each iteration $t$. Note that when the $\ell_{\infty}$-norm is used in (9), the problem can be written as an LP with $4N_{g}P+2N_{g}$ decision variables (including slack variables) and $4N_{g}P+7N_{g}$ constraints, where $N_{g}$ is the number of training samples at the $g^{th}$ client. If this problem is solved via the barrier method equipped with the log barrier function and Newton updates, it would require $\mathcal{O}(\sqrt{(}C)\log(\epsilon_{2}^{-1}))$ iterations to reach an $\epsilon_{2}$-solution [43], where $C$ is the number of constraints. Moreover, each iteration would have an arithmetic complexity of $\mathcal{O}(CD^{2})$, where $D$ is the number of decision variables. Therefore, solving the problem in (9) via the previously stated method has a theoretical worst-case time complexity of
@@
By eliminating scalar multipliers and constants, this result can simplified to the following expression,
@@
Note that all clients can solve their local problems in parallel and will have the same number of features. Thus, the client with the largest number of samples $N_{g^{\ast}}$ will have the highest time complexity. Furthermore, note that the central server performs a summation of $G+1$ vectors of dimension $P$ during each iteration $t$, the time complexity of which is $\mathcal{O}(GP)$. We obtain the final result by noting that the subgradient method converges to a solution with tolerance $\epsilon_{1}$ in $\mathcal{O}(\epsilon_{1}^{-1})$ iterations [44]. Note that we do not explicitly consider the time complexity of computing the local subgradient at each client since it is lower than that of solving the problem in (9). ∎

### v2

@@ Abstract
The training of classification models for fault diagnosis tasks using geographically dispersed data is a crucial task for original equipment manufacturers (OEMs) seeking to provide long-term service contracts (LTSCs) to their customers. Due to privacy and bandwidth constraints, such models must be trained in a federated fashion. Moreover, due to harsh industrial settings the data often suffers from feature and label uncertainty. Therefore, we study the problem of training a distributionally robust (DR) support vector machine (SVM) in a federated fashion over a network comprised of a central server and $G$ clients without sharing data. We consider the setting where the local data of each client $g$ is sampled from a unique true distribution $\mathbb{P}_{g}$, and the clients can only communicate with the central server. We propose a novel Mixture of Wasserstein Balls (MoWB) ambiguity set that relies on local Wasserstein balls centered at the empirical distribution of the data at each client. We study theoretical aspects of the proposed ambiguity set, deriving its out-of-sample performance guarantees and demonstrating that it naturally allows for the separability of the DR problem. Subsequently, we propose two distributed optimization algorithms for training the global federated distributionally robust support vector machine (FDR-SVM): i) a subgradient method-based algorithm, and ii) an alternating direction method of multipliers (ADMM)-based algorithm. We derive the optimization problems to be solved by each client and provide closed-form expressions for the computations performed by the central server during each iteration for both algorithms. Finally, we thoroughly examine the performance of the proposed algorithms in a series of numerical experiments utilizing both simulation data and popular real-world datasets.
@@ 1 Introduction
Original Equipment Manufacturers (OEMs) of capital-intensive assets such as gas turbines, aircraft engines, and mining equipment are increasingly implementing long-term service contracts (LTSCs) that can last for 10 to 20 years to ensure their products perform reliably over time [1]. Such contracts are beneficial to both the OEMs and the customers, as they provide the OEMs with a steady revenue stream, whereas they relieve the customers from all diagnosis and repair tasks in the occurrence of fault events. Indeed, one of the key services provided by the OEM as part of LTSCs is the diagnosis of any faults that may occur in the equipment used by the customers. Recently, many research efforts have been dedicated to performing fault diagnosis tasks by leveraging a wide array of data-driven Machine Learning (ML) classifiers. For example, ML models have been used in the fault diagnosis of motors [2], electrical components [3], and pumps [4].
@@
    We propose the MoWB ambiguity set, which extends the classical Wasserstein ball to the distributed setting. Our proposed ambiguity set contains mixtures of distributions from Wasserstein balls defined locally at each client. Using this ambiguity set, we reformulate the FDR-SVM problem into a separable form that can be solved in a distributed manner without sharing data. We also derive out-of-sample performance guarantees for the MoWB ambiguity set under certain assumptions.
@@ 2.1 Distributionally Robust Classifiers
Many existing works in the literature seek to develop DR versions of common classifiers to improve their out-of-sample performance. For example, the authors in [7] and [15] develop Wasserstein DR versions of logistic regression and hard and soft-margin SVMs, respectively. They provide tractable, convex reformulations for their proposed DR problems, leveraging popular duality results from [10, 9]. This work is extended in [23], where the authors derive a tractable, convex reformulation of a Wasserstein DR logistic regression model for data with mixed features, demonstrating that considering categorical data leads to a fundamentally different problem. Another extension is presented in [24], where the authors derive a DR version of the linear and kernel multiclass SVMs characterized by the Crammer-Singer loss, as well as a scalable algorithm to train a specific case of the linear model.
@@
Finally, the authors in [20] utilize group DRO to address the tendency of classification deep neural networks (DNNs) to learn spurious correlations in the data. More specifically, they manually group the training data according to those correlations and minimize the risk with respect to the worst-case group. However, they key challenge in that work is that data grouping has to be performed manually, which can be extremely difficult for very large datasets. This challenge is addressed by the authors in [26], where they utilize a DNN to perform the data grouping. The common aspect shared between the previously listed works is that they rely on the training data being present at a central location. Indeed, as argued in [18] this can make such models difficult to extend to distributed settings.
@@ 2.3 Distributionally Robust Federated Learning
Another research direction that combines elements of DRO and FL is explored in [38] where the authors develop a Wasserstein ambiguity set using barycenters. In this setting, the authors consider multiple clients, each of which has local samples from $\mathbb{P}$ allowing it to construct a local empirical distribution ^$\widehat{\mathbb{P}}_{N_{g}}$. The global ambiguity set is then centered around the barycenter (which is similar to a mean) of all {^$\{\widehat{\mathbb{P}}_{N_{g}}\}_{g=1}^{G}$. However, the barycenter is not always guaranteed to exist and can be difficult to compute in a distributed fashion when it does. Similarly, the authors in [18] study a multi-client network where each client has access to a finite number of samples from $\mathbb{P}$, and the clients seek to collaboratively solve a Wasserstein DRO problem. The authors show that the reformulation of the Wasserstein DRO problem in [10, 9] admits a reformulation that can be solved in a distributed fashion via Lagrange duality if the clients can communicate with each other. Similarly, the authors in [39] utilize Wasserstein DRO to train a DR model in a federated fashion while only allowing server to client communication. However, in proving the convergence of their proposed algorithm they assume the Lipschitz smoothness and strong convexity of the loss function under study, making the algorithm unusable for many prominent ML models such as the SVM. Moreover, they assume that the distance between any local gradient and the global gradient is uniformly bounded from above, which can be difficult to guarantee in a real-world setting.
@@ 4.1 Problem Separability
In this section, we extend the classical Wasserstein amiguity set to the distributed setting via the novel MoWB ambiguity set $\mathcal{A}_{G}$ defined next.
@@
Observe that when $G=N$, our ambiguity set models worst-case perturbations in individual training samples, rendering it a sample-driven approach with similarities to robust optimization (RO). Alternatively, when $G=1$, our proposed ambiguity set reduces to the classical Wasserstein ball $\mathcal{A}_{\varepsilon,q,d}(\Xi)$ defined in (2). This suggests that our proposed ambiguity set offers more flexibility in modeling the uncertainty than the classical Wasserstein ambiguity set, which can allow it to achieve improved performance in some settings. This also suggests that our proposed ambiguity set naturally extends the classical Wasserstein ball to the distributed setting, allowing for the problem to be reformulated in a separable fashion as we demonstrate next.
@@
$\inf_{\boldsymbol{w}}\sup_{\mathbb{Q}\in\mathcal{A}_{G}}\mathbb{E}^{\mathbb{Q}}[\ell_{H}(\boldsymbol{w};\boldsymbol{\xi})]$ $=\inf_{\boldsymbol{w}}\sup_{\left\{\mathbb{Q}_{g}\in\mathcal{A}^{(g)}_{\varepsilon_{g},1,d}(\Xi)\right\}_{g=1}^{G}}\mathbb{E}^{\sum_{g=1}^{G}\alpha_{g}\mathbb{Q}_{g}}[\ell_{H}(\boldsymbol{w};\boldsymbol{\xi})]$ (6a)
$=\inf_{\boldsymbol{w}}\sup_{\left\{\mathbb{Q}_{g}\in\mathcal{A}^{(g)}_{\varepsilon_{g},1,d}(\Xi)\right\}_{g=1}^{G}}\sum_{g=1}^{G}\alpha_{g}\mathbb{E}^{\mathbb{Q}_{g}}[\ell_{H}(\boldsymbol{w};\boldsymbol{\xi})]$ (6b)
$=\inf_{\boldsymbol{w}}\sum_{g=1}^{G}\alpha_{g}\sup_{\mathbb{Q}_{g}\in\mathcal{A}^{(g)}_{\varepsilon_{g},1,d}(\Xi)}\mathbb{E}^{\mathbb{Q}_{g}}[\ell_{H}(\boldsymbol{w};\boldsymbol{\xi})],$ (6c)
where (6a) follows from the definition of the global ambiguity $\mathcal{A}_{G}$ set in (4), (6b) follows from the law of expectation over mixture distributions, and (6c) follows by recognizing that the maximization problems are separable due to each decision variable only affecting its corresponding term. ∎
@@ 4.2 Out-of-Sample Performance Guarantees
All the training samples $\{(\widehat{\boldsymbol{x}}_{(n_{g})},\widehat{y}_{(n_{g})})\}_{n_{g}=1}^{N_{g}}$ at client $g$ are IID samples from its true distribution $\mathbb{P}_{g}$.
@@
Beyond its desirable out-of-sample performance properties, our proposed MoWB ambiguity set also inherits other desirable properties from the classical Wasserstein ambiguity set listed in [9]. One key property is the ability of the MoWB ambiguity set to anticipate black swans by possibly assigning probability mass to any point in the support set $\Xi$ regardless of the structure of ^$\widehat{\mathbb{P}}_{N_{g}}$ for all $g\in[G]$.
@@ 5.1 SM Algorithm
[…]
@@
The computed local subgradients are sent to the central server, which computes a weighted sum of the subgradients using weights $\alpha_{g}$, and performs a subgradient step on the model parameters $\boldsymbol{w}$. The central server then sends the updated $\boldsymbol{w}$ to the clients and the process repeats for $T$ iterations. We write our SM algorithm explicitly as follows:
@@ 5.2 ADMM Algorithm
The ADMM algorithm requires each client $g$ to solve their own local problem and shares their optimal local model $\boldsymbol{w}^{\ast}_{g}$ with the central server. The central server then aggregates the local models to obtain an optimal global model $\boldsymbol{w}^{\ast}$, which it sends to the clients. This process repeats for $T$ iterations.
@@
$=\inf_{\boldsymbol{w}_{g},\lambda_{g}\geq 0}\lambda_{g}\varepsilon_{g}+\frac{1}{N_{g}}\sum_{n_{g}=1}^{N_{g}}\sup_{\boldsymbol{\xi}\in\Xi}\left\{\ell_{H}(\boldsymbol{w}_{g};\boldsymbol{\xi})-\lambda_{g}d(\boldsymbol{\xi},\widehat{\boldsymbol{\xi}}_{n_{g}})\right\}+\frac{\rho}{2}||\boldsymbol{w}_{g}-\boldsymbol{w}+\boldsymbol{\mu}_{g}||_{2}^{2}$ (12c)
@@ 5.3.1 SM Algorithm
The SM Algorithm 1 converges to an optimal solution $\boldsymbol{w}^{\ast}$ of the problem in (5) within a specified tolerance $\epsilon_{1}$, provided the step-size $\gamma(t)\rightarrow 0$ as $t\rightarrow\infty$ and $\sum_{t=1}^{\infty}\gamma(t)=\infty$.
@@
As established in [43], the subgradient method guarantees convergence if three conditions are met: i) the objective function is convex, ii) the objective function is Lipschitz continuous, and iii) the step-size diminishes at an appropriate rate as stated in the Theorem statement. We verify the first two conditions in Lemmas 2 and 3, whereas the third condition can be ensured by by selecting an appropriate step-size sequence. ∎
@@
Let the $\ell_{\infty}$-norm be used in the worst-case distribution problem (9), and assume that the barrier method equipped with the log barrier function and Newton updates is used to solve it. The SM algorithm 1 with an appropriate diminishing step-size [43] has a worst-case time complexity of
$\mathcal{O}\left(\epsilon_{1}^{-2}\left[N_{g^{\ast}}^{3.5}P^{3.5}\log(\epsilon_{2}^{-1})+GP\right]\right)$
where $\epsilon_{1}$ and $\epsilon_{2}$ are the tolerances of the solutions of the subgradient method and the worst-case distribution problem (9), respectively, and $N_{g^{\ast}}$ is the greatest number of samples available at any one client.
This analysis can be directly extended to cases where other norms are used in (9).
@@
We first examine the time complexity of problem (9) that each client $g$ solves at each iteration $t$. When the $\ell_{\infty}$-norm is used in (9), the problem becomes a Linear Program (LP) with $4N_{g}P+2N_{g}$ decision variables (including slack variables) and $4N_{g}P+7N_{g}$ constraints, where $N_{g}$ is the number of training samples at the $g^{th}$ client. Solving the problem via the barrier method with the log barrier function and Newton updates requires $\mathcal{O}(\sqrt{(}C)\log(\epsilon_{2}^{-1}))$ iterations to reach an $\epsilon_{2}$-solution [44], where $C$ is the number of constraints. Moreover, each iteration has an arithmetic complexity of $\mathcal{O}(CD^{2})$, where $D$ is the number of decision variables. Therefore, the theoretical worst-case time complexity of solving the problem in (9) is:
@@
By eliminating scalar multipliers and constants, we arrive at the following simplified expression,
@@
Since all clients can solve their local problems in parallel, and will have the same number of features. Thus, the client with the largest number of samples $N_{g^{\ast}}$ will have the highest time complexity. Furthermore, the central server performs a summation of $G+1$ vectors of dimension $P$ during each iteration $t$, the time complexity of which is $\mathcal{O}(GP)$. We obtain the final result by noting that the subgradient method converges to a solution with tolerance $\epsilon_{1}$ in $\mathcal{O}(\epsilon_{1}^{-1})$ iterations [45]. Note that we do not explicitly consider the time complexity of computing the local subgradient at each client since it is lower than that of solving the problem in (9). ∎